a) & b) PLEASE
Machines A and B are mutually exclusive and are expected to produce the following real cash flows. The real opportunity cost of capital is 10% pa. 7. Cash flows (S thousands) CO -100+ 120 C3 C2 +121 +121 C1 Machine +133 +110 a. Calculate the NPV of each machine. (Correct to the neatest cent.) (4 marks) b. Which machine should you buy? Why? (1 mark)
Answer #1
a) NPV of Machine A = -100+110/1.1+121/1.1^2 = $     100.00
NPV of Machine B = -120+110/1.1+121/1.1^2+133/1.1^3 = $     179.92
b) Machine B should be bought, as it has higher NPV.